OpenStudy (anonymous):

Can some one explain how to do this? 2x-12=14

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

1) add 12 to both sides 2) divide both sides by 2

OpenStudy (anonymous):

me, i like mopars

OpenStudy (anonymous):

2x-12 = 14 first add 12 to both sides 2x-12+12 = 14+12 2x = 26 Divide both sides by 2 26/2 = 13 x=13

OpenStudy (anonymous):

and its 13...

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

Welcome (: Hope that helped